# Don’t Retrieve, Navigate: Distilling Enterprise Knowledge into Navigable Agent Skills for QA and RAG

### 29 Apr 2026 v2

@@ 3. Scroll to Currency and select your
  6221 -> 1513 -> 107 -> 18 -> 1 (programs)
  6221 -> 1513 ->  68 -> 18 -> 1 (billing)
This trace demonstrates cross-branch navigation: the agent first explored the online-programs subgroup to understand the relationship between courses and currency, then navigated to the billing-documents subgroup to find the authoritative currency configuration steps. The two branches within the same skill provided complementary information that neither alone would have answered the query fully.
@@ Appendix F Positioning Comparison
Table 8 highlights three axes on which Corpus2Skill differs from prior work.
@@
Scatter/Gather relies on human users to iteratively select clusters; RAPTOR and GraphRAG delegate navigation to fixed algorithms (vector similarity or graph traversal). Corpus2Skill is unique in delegating navigation to an LLM agent that can reason about which branch to explore, backtrack when a path is unproductive, and combine evidence across branches. This agent-driven navigation is what enables the trial-and-error behavior illustrated in Figures 4 and 5.
@@
All prior systems require dedicated infrastructure at query time: a cluster index, a vector database, or a graph database. Corpus2Skill eliminates this dependency entirely—the only runtime component is the LLM itself, which receives pre-compiled skill files via the API. This “LLM only” property simplifies deployment but shifts the cost burden to input tokens, as discussed in Section 5.
@@ Appendix G Failure Analysis
Navigation miss (38 queries).
The agent retrieves documents but none match the gold article, indicating it descended into the wrong branch of the hierarchy. For example, a query about reordering CMS collection items is routed to a general site-editing skill instead of the CMS-specific one; a query about blocking time slots in a staff schedule navigates to a generic booking group rather than the staff-management leaf. These misroutes occur at the top level of the hierarchy, where broad skill summaries fail to distinguish closely related topics.
Partial navigation (19 queries).
The agent’s retrieved set includes the gold document, yet scores remain low. In these cases the gold article is retrieved alongside several non-relevant documents, diluting the context provided to the answer synthesis step. The agent correctly identified the general topic area but over-retrieved from neighboring leaf groups.
Synthesis error (3 queries).
@@
No documents retrieved (2 queries).
Two queries (“How do I clear my browser cache?” and a Gmail password reset question) resulted in zero retrieved documents. Both have gold articles in the corpus (Wix publishes help articles for these topics), but the agent could not locate a matching skill branch and correctly declined to answer rather than hallucinating.
@@
The dominant bottleneck is initial topic routing: 38 of 62 failures stem from the agent choosing the wrong top-level branch. The compact-mode ablation (Section 4.5), which produces narrower top-level clusters, partially mitigates this by providing finer-grained topic separation at the first navigation step.
@@ Appendix I Evaluation Metrics
Lexical metrics.
Token F1 tokenizes both the predicted and gold answers into word tokens and computes the harmonic mean of token-level precision and recall. BLEU (Papineni et al., 2002) measures modified n-gram precision (up to 4-grams) with a brevity penalty to penalize overly short outputs. ROUGE-1 and ROUGE-2 (Lin, 2004) measure unigram and bigram recall respectively, capturing how much of the gold answer’s content appears in the prediction.
LLM-judged metrics.
Both Factuality and Context Recall use an LLM judge (Zheng et al., 2023) following the RAGAS framework (Es et al., 2024), scoring on a 1–5 scale normalized to 0–1. Factuality evaluates whether the claims in the generated answer are factually consistent with the gold answer, independent of the retrieved context. Context Recall evaluates whether the retrieved context contains sufficient information to answer the question, by checking how many essential claims in the gold answer can be attributed to the provided context.
@@
Input token count is the total number of tokens sent to the LLM across all turns. Per-query cost is computed using the model’s published per-token pricing (input and output tokens combined). For single-shot methods, cost reflects one LLM call; for multi-turn methods (Agentic, Corpus2Skill), it reflects all turns including tool-call reasoning.

### 15 May 2026 v3

@@ 3. Scroll to Currency and select your
  6221 -> 1797 -> 119 ->  9 -> 1 (programs)
  6221 -> 1682 -> 268 -> 12 -> 1 (currency)
This trace demonstrates cross-branch navigation: the agent first explored the online-programs branch in one skill to understand the relationship between courses and currency, then cross-jumped to the payment-provider branch in a different skill to find the authoritative currency configuration steps. The two branches across two skills provided complementary information that neither alone would have answered the query fully.
@@ Appendix F Positioning Comparison
Table˜5 highlights three axes on which Corpus2Skill differs from prior work.
@@
Scatter/Gather relies on human users to iteratively select clusters; RAPTOR and GraphRAG delegate navigation to fixed algorithms (vector similarity or graph traversal). Corpus2Skill is unique in delegating navigation to an LLM agent that can reason about which branch to explore, backtrack when a path is unproductive, and combine evidence across branches. This agent-driven navigation is what enables the trial-and-error behavior illustrated in Appendix˜E.
@@
All prior systems require dedicated infrastructure at query time: a cluster index, a vector database, or a graph database. Corpus2Skill eliminates this dependency entirely—the only runtime component is the LLM itself, which receives pre-compiled skill files via the API. This “LLM only” property simplifies deployment but shifts the cost burden to input tokens, as discussed in Limitations.
@@ Appendix G Failure Analysis
Navigation miss (19 queries).
The agent retrieves documents but none overlap with the gold context (Context Recall==0), indicating it descended into the wrong branch. For example, a query about connecting PayPal is routed to a general payments group instead of the payment-setup leaf; a query about changing a header background on scroll navigates to a layout group rather than the header-effects cluster. These misroutes typically occur at the top level, where broad skill summaries fail to distinguish closely related topics.
Partial navigation (16 queries).
The agent’s retrieved set partially overlaps with the gold (0<0<Context Recall<0.6<0.6) but includes too many non-relevant documents, diluting the context and producing a low-quality answer. The agent identified the correct general topic area but over-retrieved from neighboring leaf groups.
Synthesis error (5 queries).
@@
No documents retrieved (5 queries).
Five queries resulted in zero or near-zero retrieval. These include off-domain queries (“How do I clear my browser cache?”, a Gmail password reset question, Google Ads billing) where the agent could not locate a matching skill branch and declined to answer rather than hallucinating.
@@
The dominant bottleneck remains initial topic routing: 19 of 45 failures stem from the agent choosing the wrong top-level branch. The narrow-tree ablation (Section˜5.4), which produces finer-grained top-level clusters, partially mitigates this by providing sharper topic separation at the first navigation step.
@@ Appendix I Evaluation Metrics
Lexical and semantic answer-quality metrics.
Token F1 tokenizes both the predicted and gold answers into word tokens and computes the harmonic mean of token-level precision and recall. BERTScore-F1 (Zhang et al., 2020) computes contextual-embedding similarity F1 between the predicted and gold answers using a RoBERTa-large encoder (Liu et al., 2019) with rescale_with_baseline=False; it is more robust to paraphrase than n-gram-overlap metrics. We compute it once per query and report the mean.
LLM-judged answer-quality and grounding metrics.
[…]
Hallucination Rate.
[…]
@@
[…]